OpenStudy (wild_child117):

The sum of three consecutive even integers is 72. Let x represent the least consecutive integer. What equation could be used to find the least consecutive integer? What is the second consecutive integer? Enter your answers in the boxes. Equation: Second consecutive integer:

OpenStudy (wild_child117):

@Awolflover1 can you please help

OpenStudy (wild_child117):

please help my grade is resting on this quiz

OpenStudy (wild_child117):

hey @Cas_fangirl_14 do you think you can help me

OpenStudy (cas_fangirl_14):

I haven't learned this yet. Sorry. :(

OpenStudy (wild_child117):

do you know anyone who does?

OpenStudy (mathmale):

The problem statement tells you to let the first integer be x. What is an algebraic expression for the next consecutive integer? of the next one after that?

OpenStudy (mathmale):

First integer: x Second int. : ? Third int. : ?

OpenStudy (mathmale):

If you're unsure, please look up the word "consecutive." Then look up the phrase "consecutive integers."

OpenStudy (wild_child117):

i looked it up but i still dont understand the problem

OpenStudy (sshayer):

suppose first 24,next 26,next28,... or 24,24+2,24+4 ,... if first x then next two are...

OpenStudy (wild_child117):

24+6?

OpenStudy (sshayer):

i asked if first x ,then next even integer is ...

OpenStudy (sshayer):

next even integer is x+2 now what is the next even integer?

OpenStudy (wild_child117):

2+4

OpenStudy (wild_child117):

im not good at this sorry

OpenStudy (sshayer):

if first 10 next even integer 10+2,and next even integer is 10+4,... similarly if first x is even integer then next even integer is x+2 and next even integer is...

OpenStudy (sshayer):

next even integer is x+4 now x+x+2+x+4=72 find x

OpenStudy (mathmale):

sshayer's example, suppose first 24,next 26,next28,... or 24,24+2,24+4 ,..., refers to CONSECUTIVE EVEN INTEGERS. 24, 26, 28, ..., are all even. In this problem you are asked to represent and evaluate 3 consecutive integers, meaning that the difference between the 1st and 2nd is 1, the diff between the 2nd and 3rd is also 1, and so on. So, if the first integer is represented by x, the next consecutive integer is .... ? ?

OpenStudy (sshayer):

i think question wants even consecutive integers.

OpenStudy (wild_child117):

22 + 24 + 28?

OpenStudy (wild_child117):

am i right @mathmale sorry to bother you again

OpenStudy (wild_child117):

22+24+26 my bad

jhonyy9 (jhonyy9):

yes you are right with these bc. 22+24+26 =72

OpenStudy (mathmale):

I am sorry that my EXAMPLE has caused confusion. Yes, you are to find 3 consecutive integers. If the first integer is x, what is the next integer? the next? (3 in all). You'll need to practice converting word problems such as this one into appropriate symbols and equations. The sum of the 3 consec. int. is 72: x + ( ) + ( ) = 72 second int. third int.